mirror sites: PL (internal link) HQ     [?]

Health Check monitor

 

CAL | HC | refs | QC
HealthCheck Monitor
   HOME | UsersGuide
  sl icon ALL INSTRUMENTS
 
FULL reports
VISIR:
detectors
Other HC:
QC links:
        [page auto-refreshes after 300 sec]   [stop | on]            [press Ctrl+Shift+R to enforce refresh of scores, dates and news]
VISIR trending system: HEALTH CHECK report
HELP USERS-GUIDE MORE
Last update: 2014-10-26T02:00:41 (UT)
  | now: 2014-10-26T08:14:47 (UT) muc03 QC pipeline: visir-3.5.0 (installed 2014-10-20)
same group: read-noise read-noise12 dark flat-flux nbadpix det-temp
General news:
VISIR news: 2012-10-18: VISIR is currently not operational
Report news:
DATE*: [?] 2014-10-19
no data | NLT
2014-10-20
no data | NLT
2014-10-21
no data | NLT
2014-10-22
no data | NLT
2014-10-23
no data | NLT
2014-10-24
no data | NLT
2014-10-25
no data | NLT
P...
Product quality:
no products no products no products no products no products no products no products
  last OPSLOG data:
2014-10-26*

last files ...
monitor

scores&comments | history ... | plot tutorial ... | contact                 daily/often; important to check     [?]
*Date on this monitor changes at 21:00 UT
Print: Information and research:
- png file - advanced studies: QC1 browser | QC1 plotter
Plot
?
Symb
?
Source
*
OPS?
**
Average ? Thresholds ? N_
data
QC1
parameter
Data
downloads
Remarks
method value unit method value
1 QC1DB yes (0) MEAN   ADU VAL45,70 0 ron2 this | last_yr | all read-out-noise of the imaging detector
2 QC1DB yes (0) MEAN   ADU VAL0,70 0 ron2 this | last_yr | all read-out-noise of the spectroscopic detector
*Data sources: QC1DB: QC1 database; LOCAL: local text file | **OPS: OPSLOG data are included
Plot 1
scores: 2|
data source:visir_dark_img
(QC1 database)
dataset:ron2
mean:no dataADU
fixed thresholds:45...70ADU
N_data plotted:0
[click on plot for closeup]
Plot 2
scores: 4|
data source:visir_dark_spc
(QC1 database)
dataset:ron2
mean:no dataADU
fixed thresholds:0...70ADU
N_data plotted:0
[click on plot for closeup]
This plot

This is a trending plot monitoring the read-out-noise (RON) measured from the raw(3)-raw(2) difference frame. It's been introduced since November 2007. The RON values are displayed for both, the imaging and spectroscopic detectors. The data points for the spectroscopic detector include those taken in MR (values around 150) and in HRS (values around 180).  


General information

Click on any of the plots to see a close-up version.

If applicable, the latest values from Paranal ops logs are plotted as blue dots. More information about those values, including a download option, is available in the OPS column.

The latest date is indicated on top of the plot, data points belonging to that date are specially marked.

If configured,

  • statistical averages are indicated by a solid line, and thresholds by broken lines
  • outliers are marked by a red asterisk. They are defined as data points outside the threshold lines
  • "aliens" (= data points outside the plot Y limits) are marked by a red arrow ( or )
  • you can download the data for each parameter set if the 'Data downloads' link shows up